Find the area of the shaded region. The graph depicts the standard normal distribution of bone density scores with mean 0 and standard deviation 1. font size decreased by 2 z equals negative 0.97font size decreased by 2 z equals 1.21
Answer #1

Answer:

Given,

Mean = 0

Standard deviation = 1

To give the area of region b/w z = - 0.97 & z = 1.21

P(- 0.97 < X < 1.21) = P((-0.97 - 0)/1 < z < (1.21 - 0)/1)

= P(-0.97 < z < 1.21)

= P(z < 1.21) - P(z < - 0.97)

= 0.8868606 -  0.1660232 [since from z table]

= 0.7208374

P(- 0.97 < X < 1.21) = 0.7208